Find the total force on the side of a water-filled tube 2.34 cm high with a radius of 0.300 cm.
1 : 0.506 N
2 : 0.0506 N
3 : 3.41 N
4 : 1.03 N
1
Expert's answer
2020-06-15T10:36:41-0400
"dF=\rho gh dA=2\pi r\rho gh d h"

"F=\pi r\rho gh ^2=\pi (1000)(9.8)(0.003)(0.0234) ^2\\=0.0506\ N"